OpenStudy (anonymous):

A triangle with a base of 1/4 meter has an area of 8 square meters. What is the height, in meters, of the triangle? a.1 b.12 c.32 d.64

OpenStudy (anonymous):

1/2*b*h=A you have the base and area (A), you want the height (h) so change the formula. h=A/(1/2*b) h=8/(1/2*1/4)